Stephanie drove at an average rate of 50 miles per hour for 2 hours and then increased her average rate by 50% for the next 3 hours. Her average rate of speed for the 5 hours was t miles per hour. What is the value of t?
Answer by Nate(3500)

average ~> 50mph
increase of 50% of average ~> 1.5(50) = 75mph
(50*2 + 75*3)/5 = t
(100 + 225)/5 = t
325/5 = t
65mph = t